3 major projects P are being funded by 3 voluntary agencies 1 P,
2 P,
3 1 2 V3 V . ,V ,
1 2 V3 V are willing to pay Rs. 8,000/-, Rs. 4,000/- and R ,V , s. 2, 000/-, respectively per
person on the project P ; Rs. 4,000/-, Rs. 3,000/- and Rs. 4, 000/- respec 1
tively per
person on P ; and Rs. 3,000/-, Rs. 5,000/-, Rs. 8,000/- respec 2
tively on the project P . 3
Further, the amount that 1 2 V3 V have kept aside for paying people on these projec ,V , ts
is Rs. 2,17,000/-, Rs. 1,42,000/- and Rs. 1,32,000/- respectively. How many people
should each project employ so that the total money available is utilised?
1
Expert's answer
2019-04-09T10:00:52-0400

x- people on P1

y-people on P2

z-people on P3


8000x+4000y+3000z=217000

4000x+3000y+5000z=142000

2000x+4000y+8000z=132000


8x+4y+3z=217

4x+3y+5z=142

2x+4y+8z=132


8x+4y+3z=217. 8x+4y+3z=217. - 12y-29z=-311

x+2y+4z=66. - 8x-16y-32z=-528


4x+3y+5z=142. 4x+3y+5z=142. - 5y-11z=-122

x+2y+4z=66. - 4x-8y-16z=-264


-12y-29z=-311. 132y+319z=3421. - 13y=-117

-5y-11z=-122. - 145y-319z=-3538


y=9

z=7

x=66-18-28=20


Answer: P1-20; P2-9; P3-7